1.A Triangle Has Vertices at A(3,2) B(4,0) C(8,2).
a) Show the triangle is right angled
b) Find the Co-ordinates of the Point D such that ABCD is a rectangle

Answer by ewatrrr(10682)


Hi,
a. Show the triangle is right angled
mAB = -2/1 mBC = 2/4 = 1/2
AB is perpendicular to BC as their slopes are negative reciprocals.
b. Co-ordinates of the Point D such that ABCD is a rectangle
BC and AD would be parallel and have same slope = 1/2
y = 1/2x + b Using point (3,2) b = -1/2
y = .5x + .5 (red line) mCD = -2 CD meets AD at (7,4)